Grade 9Mechanics

A​ particle is thrown with a velocity u. At t=2 s the particle makes an angle 30 degree with the horizontal. At t=3 s it reaches the maximum height. Find the angle of projection and u.

Sayak
10 Years agoGrade 9

1 Answer

arun
10 Years ago
let initial velocity of projectile be ‘u‘. And angle of projection be \theta 
so initial velocity in horizontal dirction is u\cos\theta
and initial velocity in vertical direction is u\sin\theta
velocity of projectile at t=2 sec is v
firstly considering the vertical motion
at t=3 sec
from first equation of motion 
0= u\sin\theta -g(3)
u=\frac{3g}{\sin\theta}                                                        …....1
 
at t= 2 sec
from first equation of motion 
v\sin30=u\sin\theta-g(2)
from equation 1
\frac{v}{2}=3g-2g
v=2g                                                    …......2
now considering the horizontal motion 
at t=2 sec
from first equation of motion
v\cos30=u\cos\theta-(0)(2)
from equation 2
u=\frac{\sqrt{3}g}{\cos\theta}
from equation 1
\frac{\sqrt{3}g}{\cos\theta}=\frac{3g}{\sin\theta}
\tan \theta=\sqrt{3}
\theta=60
from equation 1
u=2\sqrt{3}g
hence u= 2(3)^(1/2)g         and angle of projection is 60